# Traffic Scenario Logic: A Spatial-Temporal Logic for Modeling and Reasoning of Urban Traffic Scenarios

### v1

@@ 6 Application
TSL is a useful tool for modeling and reasoning traffic scenarios. It can generate various traffic scenarios, which can be used to create a library of test cases for comprehensive testing of autonomous vehicles. (Zhang et al., 2021; Bagschik, Menzel, and Maurer, 2018) Additionally, TSL can identify potential failure risks during the development of autonomous driving strategies, which is our previous work. (Wang et al., 2023) Its ability to differentiate between homotopy classes of scenarios makes it well-suited for these applications, which require both the comprehensiveness and generalization of scenarios.
In this section, we start by introducing the reasoning of TSL, which we implemented using answer set programming with temporal logic extensions. Following that, we illustrate several examples of TSL for scenario generation to demonstrate its effectiveness in handling miscellaneous urban scenarios.
@@ 6.1 Reasoning of TSL
Given an initial scene $\mathfrak{s}_{0}$, the set of vehicles $\mathbb{C}$, a finite horizon $T$ and a road network $\mathcal{M}$, the reasoning of TSL is to find all logical scenarios $\mathcal{S}=(\mathbb{C},\mathbb{L},\mathbb{D},\mathbb{P},T,\langle\mathcal{S}_{0},\mathcal{S}_{1},\dots,\mathcal{S}_{T-1}\rangle)$ that $\mathcal{M},\mathcal{S}\models\Phi$, where $\Phi$ is the set of all rules explained in Section 4 and Section 5.
We use Telingo (Cabalar et al., 2019) to implement the reasoning of TSL. Telingo is a solver for temporal logic programs based on the Answer Set Programming (ASP) and the well-known Clingo (Gebser et al., 2019) input language. Generally, a logic program of the Answer Set Programming consists of rules like
A0 :- L1,L2,…,Ln.
where the head A0 is an atom, and the body is the conjunction of literals Li$(1\leq i\leq n)$ of the form A or not A, where A is an atom and not represents the Negation as Failure (NAF). A rule without a body is a fact, and a rule without a head becomes a constraint. Telingo extends the ASP language by importing temporal modal operators under the semantics of the Linear Temporal Logic ($\mathcal{LTL}$(Cabalar et al., 2018; Pnueli, 1977).
We use aggregates (Son and Pontelli, 2007) in the logic program we used for TSL reasoning to derive different scenarios. All rules introduced in Section 4 and Section 5 are rewritten into corresponding ASP rules or constraints. The road network and initial scene are described by facts. For the entire implementation of our reasoning programs, please refer to our source code [^2] https://github.com/Autonomous-Driving-Safety-Project/TSL.
@@ 6.2 Examples of Scenario Generation
In this subsection, we provide several examples of TSL on different road layouts.
Example 1 (Highway scenario)
4(a) shows a uni-direction two-lane highway. The road network can be written as:
#program always.
is_road(r1).
is_lane(l1;l2).
has_lane(r1,l1).
has_lane(r1,l2).
left(l1,l2).
The initial scene is:
#program initial.
on_lane(c1,l2).
on_lane(c2,l2).
lonr(c1,c2,behind).
By adding a constraint that c2 shall not be in front of c1 in the final scene:
#program final.
:- lonr(c2,c1,ahead).
We can obtain the 4 shortest overtaking scenarios in 3 steps, where the first one is
State 0:
lonr(c2,c1,ahead) lonr(c1,c2,behind)
on(c1,l2) on(c2,l2)
State 1:
lonr(c2,c1,ahead) lonr(c1,c2,behind)
on(c1,l1) on(c1,l2) on(c2,l2)
State 2:
lonr(c1,c2,cover) lonr(c2,c1,cover)
on(c1,l1) on(c2,l2)
All models we get are shown in 4(a). In the first scenario, c1 (the blue one) overtakes c2 (the red one) by lane change. Two cars change lanes simultaneously in the second and third scenarios, and then one gives up. In the fourth scenario, c2 yields to c1.
@@
When two cars pass an intersection point, TSL gives two different scenarios, as shown in 4(b). Please note that the two lanes are not connected at the intersection point, so a car cannot switch to another lane.
@@
4(c) illustrates a scenario where a car passes through a connection point and may drive into any of the connecting lanes.
@@
When a vehicle passes through two intersection points successively, there are two possible scenarios since the length of the vehicle and the distance between the intersection points are unknown, as shown in 4(d). A long truck may cover both intersection points simultaneously, while a short car can be in between.
@@
TSL can use overlapping to create overtaking on opposite scenarios, as shown in 4(e). It contains two roads: one road has two lanes, l1 and l2, running from left to right, and the other road has one lane, l3, running from right to left, with l3 overlapping with l2. Vehicles c1 (blue) and c2 (red) are initially located on l1, and c3 (yellow) is located on l3. With no lane changes allowed for c2 and c3, two scenarios are provided by TSL for c1 to overtake c2: waiting for c3 to pass before overtaking, and overtaking before c3 passes.
Example 6 (The T-intersection)
Figure 2 depicts a typical T-intersection, which combines the primitives shown in the previous examples. TSL reasoning still applies, but there are too many possible scenarios to list here. For the complete example, please refer to our source code.
The examples above show how TSL works for scenario generation. By combining these primitives, TSL can manage different urban road networks and create a variety of traffic scenarios. Please refer to our source code for the detailed implementation of these examples.
The scenarios generated can be described using a scenario description language and simulated as test cases for autonomous vehicles. For instance, Example 1 can serve as a test case for the overtaking maneuver of an autonomous driving system when the system takes control of c1. By examining the system’s behavior in various scenarios, we can assess whether it might fail to respond appropriately to c2. We have demonstrated the effectiveness of this workflow in highway scenarios in our previous study (Wang et al., 2023), and TSL assists in its expansion to general urban scenarios.
In addition to scenario generation, TSL can be used in other ways as an effective reasoning tool for traffic scenarios. We have developed a framework for highway decision-making of autonomous vehicles using a modified version of TSL [^3] This work is under submission to ITSC 2024. Link of the project: https://github.com/Autonomous-Driving-Safety-Project/TSL_Driver. Also, by formally modeling the control of an autonomous vehicle, we can verify the safety of control sequences using TSL, which will be one of our future works.
@@ 7 Conclusion
In this paper, we propose a novel spatial-temporal logic, Traffic Scenario Logic, for modeling and reasoning of urban pedestrian-free traffic scenarios. The compatibility with OpenDRIVE, the de facto industry standard of high-definition maps, and the capability to distinguish different homotopy classes of scenarios set TSL apart from existing approaches. These features make TSL well-suited for developing a complete and human-interpretable test scenario library, which is urgently needed by scenario-based autonomous vehicle safety assessment. We provide examples to illustrate TSL’s effectiveness in generating scenarios on various road networks and its capability in generating comprehensive scenarios under complex urban road networks. Our future work will involve incorporating pedestrians into scenarios and exploring additional applications of TSL.

### v3

@@ 5 Application
TSL is a useful tool for modeling and reasoning about traffic scenarios. It can generate various traffic scenarios, which can be used to create a library of test cases for comprehensive testing of autonomous vehicles. Additionally, TSL can identify potential failure risks while developing autonomous driving strategies. Its unique spatial representation makes it well-suited for these applications, which require both the comprehensiveness and generalization of scenarios.
In this section, we start by introducing the reasoning of TSL, which we implemented using Answer Set Programming (ASP) with temporal logic extensions. Following that, we illustrate several examples of TSL for scenario generation to demonstrate its effectiveness in handling miscellaneous urban scenarios.
@@ 5.1 Reasoning of TSL
Given an initial scene $\mathfrak{s}_{0}$, the set of vehicles $\mathbb{C}$, a finite horizon $T$ and a road network $\mathcal{M}$, the reasoning of TSL is a kind of model expansion, i.e. to find all logical scenarios $\mathcal{S}=(\mathbb{C},\mathbb{L},\mathbb{D},\mathbb{P},T,\langle\mathcal{S}_{0},\mathcal{S}_{1},\dots,\mathcal{S}_{T-1}\rangle)$ such that $\mathcal{M},\mathcal{S}\models\Phi$, where $\Phi$ is the set of rules explained in Section 3 and Section 4.
We use Telingo (Cabalar et al. 2019) to implement the reasoning of TSL. Telingo is a solver for temporal logic programs based on ASP and the well-known Clingo (Gebser et al. 2019) input language. Telingo extends the ASP language by importing temporal modal operators under the semantics of the Linear Temporal Logic ($\mathcal{LTL}$(Cabalar et al. 2018; Pnueli 1977).
We use aggregates (Son and Pontelli 2007) instead of actions and transitions in the logic program we used for TSL reasoning to derive different scenarios. All rules introduced in Section 3 and Section 4 are rewritten into corresponding ASP rules or constraints. The road network and initial scene are described by facts. For the entire implementation of our reasoning programs, please refer to our source code.
@@ 5.2 Examples of Scenario Generation
In this subsection, we provide several examples of TSL on different road layouts to provide an overview of TSL’s features in scenario reasoning.
Example 1 (Highway overtake)
Figure 2(a) shows overtake scenarios on a uni-direction two-lane highway. The shortest scenarios we get are shown in Figure 2(a). In the first scenario, c1 (the blue one) overtakes c2 (the red one) by lane change. In the second and third scenarios, the two cars change lanes simultaneously, and then one of the cars gives up midway through the lane change. In the fourth scenario, c2 changes lanes to the left to make way for c1.
@@
When two cars pass an intersection point, TSL gives two different scenarios, as shown in Figure 2(b). Please note that the two lanes are not connected at the intersection point, so a car cannot switch to the other lane.
@@
Figure 2(c) illustrates a scenario where a car passes through a connection point and may drive into any of the connecting lanes.
@@
When a vehicle passes through two intersection points successively, there are two possible scenarios since the length of the vehicle and the distance between the intersection points are unknown, as shown in Figure 2(d). A long truck may cover both intersection points simultaneously, while a short car can be in between.
@@
TSL can use overlap to create overtaking on opposite scenarios, as shown in Figure 2(e). It contains two roads: one road has two lanes, l1 and l2, running from left to right, and the other road has one lane, l3, running from right to left, with l3 overlapping with l2. With no lane changes allowed for c2 and c3, two scenarios are provided by TSL for c1 to overtake c2: waiting for c3 to pass before overtaking, or overtaking before c3 passes.
The examples above show how TSL works for scenario generation. By combining these primitives, TSL can manage different urban road networks and create a variety of traffic scenarios. Please refer to our source code for the detailed implementation of these examples. Theorem 2 shows that TSL is not limited to such simple scenarios, but can theoretically handle arbitrarily complex scenarios. More complex examples are provided in the Technical Appendix.
5.3 Use TSL for Test Scenario Generation
The scenarios generated can be described using a scenario description language and simulated as test cases for autonomous vehicles. We use OpenSCENARIO DSL (OSC) [^1] https://www.asam.net/standards/detail/openscenario-dsl/ as an example. It is a widely accepted industry standard proposed by ASAM for describing abstract scenarios, particularly for autonomous systems testing uses. For the syntax and semantics of OSC, please refer to its document.
A TSL scenario can be easily translated into OSC with the following steps (OSC keywords shown in typewriter font):
  1. 1.
    Create a top-level scenario;
  2. 2.
    Declare all points introduced in Section 4.1 with the position_3d type;
  3. 3.
    Add a do serial block to the top-level scenario. Then, for each scene in the TSL scenario, add a corresponding parallel sub-block;
  4. 4.
    In each parallel sub-block, describe the drive() action of every vehicle in this scene. The OSC provides the position modifier for the longitudinal position (relative to another vehicle or point declared in step 2) and lateral for the lateral position (relative to lane boundaries).
This translation is straightforward and can be automatized. A complete example is provided in the Technical Appendix. An OSC file can be further evaluated by a simulator to serve as a test case (Zhang et al. 2021; Bagschik, Menzel, and Maurer 2018) or used to find possible malfunctions (Wang et al. 2023). Specific usages of OSC files are beyond the scope of this paper.
5.4 Discussion
TSL has numerous potential applications beyond test scenario generation. For example, it can accomplish autonomous driving decision-making in a more interpretable manner and seamlessly integrate predictions for other vehicles into its reasoning process. A demonstration is provided in the Technical Appendix. It can also be used to verify the safety of control sequences for autonomous vehicles, which is a focus of our future research.
New predicates and constraints can also be introduced to extend the semantics of TSL for broader applications. For example, predicates representing vehicle types can be added, and additional traffic rules specific to certain vehicle types can be incorporated.
@@ 6 Conclusion
In this paper, we propose a novel spatial-temporal logic, Traffic Scenario Logic, for modeling and reasoning of urban pedestrian-free traffic scenarios. TSL is compatible with OpenDRIVE and has an efficient and complete spatial representation, making it well-suited for developing a human-interpretable test scenario library for autonomous vehicle safety assessment. Our future work will involve incorporating pedestrians into scenarios and exploring additional applications of TSL.
Acknowledgments
This work was partially supported by National Key R&D Program of China No. 2023YFB4704500, Hunan Province Major Scientific and Technological Project No. 2024QK2001, and was partially supported by Sangfor Technologies Co., Ltd.